A train stopping at 2 stations 4 km apart takes 4 min on the journey from one of the station to the other.Assuming that it first accelerates with uni.acceleration x and then that of uniform retardation y, prove that (1/x)+(1/y)=2

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

let t1 is time of acceleration and s1 is distance covered in this time .v is the velocity after time t

let t2 is time of retardation and s2 is distance covered in this time .

v2=02-2xs1

s1=v2/2x..............1

v=0+xt1

v/x=t1 ..................2

 

and  0=v-yt2

   t2=v/y .................3

02=v2-2ys2

s2=v2/2y..............4

add 2 and 3

1/x  +1/y  =(t1+t2)/v

               =4/v..........5

add 1 and 4

s1+s2 =v2/2(1/x +1/y)

4=v2/2(1/x +1/y)

or (1/x +1/y) =8/v2

put value of v from equation 5

1/x +1/y  =2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin